# Reinforcement Learning via Self-Distillation

### v1

@@ B.2 Trust-region Teacher
The solution to Equation˜13 can be expressed in closed form as
@@
Let $\alpha:=1/\lambda$. Then, the solution to Equation˜13 can be characterized in closed form as
@@
B.3 EMA Teacher as an Implicit Trust Region
To stabilize training, an alternative to the explicit trust-region teacher in Appendix B.2 is to parameterize the teacher as an exponential moving average (EMA) of the student’s parameters $\theta_{k}$:
$\theta^{\prime}_{k}=(1-\alpha)\theta^{\prime}_{k-1}+\alpha\theta_{k},\qquad\alpha\in(0,1),$ (15)
with initialization $\theta_{0}=\theta^{\prime}_{0}={\theta_{{\mathrm{ref}}}}$. While Equation˜15 constrains the teacher in parameter space (and thus does not, in general, impose an explicit KL trust region in distribution space), under a mild smoothness condition we can show that the EMA teacher $q_{\theta^{\prime}}$ remains within an approximate trust region relative to the reference teacher $q_{\theta_{\mathrm{ref}}}$.
Let $q_{\theta}(y_{t}\mid x,f,y_{<t})$ be a softmax distribution with logits $z_{\theta}(x,f,y_{<t})\in\mathbb{R}^{|\mathcal{V}|}$:
$q_{\theta}(y_{t}\mid x,f,y_{<t})=\mathrm{softmax}\big(z_{\theta}(x,f,y_{<t})\big)_{y_{t}}.$ (16)
Assume the logits are $L$-Lipschitz in parameters:
$\big\|z_{\theta_{1}}(x,f,y_{<t})-z_{\theta_{2}}(x,f,y_{<t})\big\|_{2}\leq L\,\big\|\theta_{1}-\theta_{2}\big\|_{2},\qquad\forall(x,f,t,\theta_{1},\theta_{2}).$ (17)
Proposition B.2 (EMA yields an approximate trust region).
Let $\theta_{0}=\theta^{\prime}_{0}={\theta_{{\mathrm{ref}}}}$, $k\geq 1$, and update $\theta^{\prime}_{k}$ by Equation˜15. We assume:
  • The logits are $L$-Lipschitz in parameters (cf. Equation˜17).
  • The student does not diverge from the initial model, i.e., $\|\theta_{i}-\theta_{0}\|_{2}\leq R_{\mathrm{ref}}$ for all $i\leq k$.
Then, for any $(x,c)$ and any sequence length $T$, the EMA teacher $\smash{q_{\theta^{\prime}_{k}}}$ implicitly satisfies the trust-region constraint
$\sum_{t=1}^{T}\mathrm{KL}\left(q_{\theta^{\prime}_{k}}(\cdot\mid x,f,y_{<t})\|q_{\theta_{\mathrm{ref}}}(\cdot\mid x,f,y_{<t})\right)\leq\epsilon_{k},\qquad\epsilon_{k}=\frac{L^{2}}{4}\,T\,R_{\mathrm{ref}}^{2}\,(1-(1-\alpha)^{k})^{2}.$ (18)
Proof.
Fix $(x,f,t)$ and define $p=\mathrm{softmax}(a)$ and $q=\mathrm{softmax}(b)$ with $a=z_{\theta^{\prime}_{k}}(x,f,y_{<t})$ and $b=z_{\theta_{\mathrm{ref}}}(x,f,y_{<t})$. A standard smoothness bound for softmax distributions [^10] Let $p=\mathrm{softmax}(a)$ and $q=\mathrm{softmax}(b)$, and define the log-partition $A(z)=\log\sum_{i}\exp(z_{i})$. For the categorical exponential family, the KL divergence admits the (primal) Bregman form $\mathrm{KL}\left(p\|q\right)=A(b)-A(a)-\langle\boldsymbol{\nabla}A(a),\,b-a\rangle$ (Wainwright & Jordan, 2008, Eq. (5.10)). Moreover, $\boldsymbol{\nabla}A(z)=\mathrm{softmax}(z)$ and $\boldsymbol{\nabla}^{2}A(z)=\mathrm{diag}(p)-pp^{\top}$ (Boyd & Vandenberghe, 2004). For any $v\in\mathbb{R}^{|\mathcal{V}|}$, $v^{\top}\boldsymbol{\nabla}^{2}A(z)v=\sum_{i}p_{i}v_{i}^{2}-\Big(\sum_{i}p_{i}v_{i}\Big)^{2}=\mathrm{Var}_{i\sim p}(v_{i})\leq\frac{(\max_{i}v_{i}-\min_{i}v_{i})^{2}}{4}\leq\frac{\|v\|_{2}^{2}}{2},$ so $\|\boldsymbol{\nabla}^{2}A(z)\|_{\mathrm{op}}\leq\tfrac{1}{2}$ and hence $A$ is $(1/2)$-smooth in $\|\cdot\|_{2}$. Applying the standard smoothness inequality for $\beta$-smooth functions (Bubeck, 2015, Lemma 3.4) yields $A(b)\leq A(a)+\langle\boldsymbol{\nabla}A(a),b-a\rangle+\tfrac{\beta}{2}\|b-a\|_{2}^{2}$; setting $\beta=\tfrac{1}{2}$ gives $\mathrm{KL}\left(p\|q\right)\leq\tfrac{1}{4}\|a-b\|_{2}^{2}$. implies
$\mathrm{KL}\left(p\|q\right)\leq\frac{1}{4}\,\|a-b\|_{2}^{2}.$ (19)
By Equation˜17, this gives
$\mathrm{KL}\left(q_{\theta^{\prime}_{k}}(\cdot\mid x,f,y_{<t})\|q_{{\theta_{{\mathrm{ref}}}}}(\cdot\mid x,f,y_{<t})\right)\leq\frac{L^{2}}{4}\,\|\theta^{\prime}_{k}-{\theta_{{\mathrm{ref}}}}\|_{2}^{2}.$ (20)
Unrolling the EMA recursion with $\theta^{\prime}_{0}={\theta_{{\mathrm{ref}}}}$ yields
$\theta^{\prime}_{k}-{\theta_{{\mathrm{ref}}}}=\sum_{i=1}^{k}w_{i}(\theta_{i}-{\theta_{{\mathrm{ref}}}}),\qquad w_{i}=\alpha(1-\alpha)^{k-i},\qquad\sum_{i=1}^{k}w_{i}=1-(1-\alpha)^{k}.$
By Cauchy–Schwarz,
$\Big\|\sum_{i=1}^{k}w_{i}(\theta_{i}-{\theta_{{\mathrm{ref}}}})\Big\|_{2}^{2}\leq\Big(\sum_{i=1}^{k}w_{i}\Big)\Big(\sum_{i=1}^{k}w_{i}\|\theta_{i}-{\theta_{{\mathrm{ref}}}}\|_{2}^{2}\Big)=(1-(1-\alpha)^{k})\sum_{i=1}^{k}w_{i}\|\theta_{i}-{\theta_{{\mathrm{ref}}}}\|_{2}^{2}.$ (21)
Combining Equations˜20 and 21 and summing over $t=1,\dots,T$ gives
$\sum_{t=1}^{T}\mathrm{KL}\left(q_{\theta^{\prime}_{k}}(\cdot\mid x,f,y_{<t})\|q_{{\theta_{{\mathrm{ref}}}}}(\cdot\mid x,f,y_{<t})\right)\leq\frac{L^{2}}{4}\,T\,(1-(1-\alpha)^{k})\,\sum_{i=1}^{k}\alpha(1-\alpha)^{k-i}\,\big\|\theta_{i}-{\theta_{{\mathrm{ref}}}}\big\|_{2}^{2}.$
Finally, if $\|\theta_{i}-{\theta_{{\mathrm{ref}}}}\|_{2}\leq R_{\mathrm{ref}}$ for all $i\leq k$, then $\sum_{i=1}^{k}w_{i}\|\theta_{i}-{\theta_{{\mathrm{ref}}}}\|_{2}^{2}\leq R_{\mathrm{ref}}^{2}\sum_{i=1}^{k}w_{i}=R_{\mathrm{ref}}^{2}(1-(1-\alpha)^{k})$, which yields Equation˜18. ∎
To summarize, Equation˜18 recovers the same form as the explicit trust-region constraint $\smash{\sum_{t}\mathrm{KL}\left(\cdot\|q_{{\theta_{{\mathrm{ref}}}}}\right)\leq\epsilon}$ (cf. Section˜B.2), with an effective radius $\epsilon_{k}$ controlled by the EMA rate $\alpha$ and the cumulative deviation from the reference. For fixed $k$ and small $\alpha$, $\smash{1-(1-\alpha)^{k}\approx\alpha k}$, so the reference divergence scales as $\epsilon_{k}=O(\alpha^{2}k^{2})$ when $\theta_{i}$ stays within a bounded neighborhood of $\theta_{{\mathrm{ref}}}$}.
@@ Appendix C Additional Related Work
Several recent works have explored assigning dense (per-token) rewards given access to an external reward model, leveraging internal structure of the reward model (Chan et al., 2024; Cao et al., 2025).
@@ C.1 SDPO as Maximum Entropy RL
where $\smash{\pi_{\theta}(y\mid x)=\prod_{t=1}^{T}\pi_{\theta}(y_{t}\mid x,y_{<t})}$ and $\smash{\mathrm{H}\left[\pi_{\theta}(\cdot\mid x)\right]=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}{}\left[-\log\pi_{\theta}(y\mid x)\right]}$ is the entropy of the policy. Here, $r(y_{t}\mid x,y_{<t})$ is an arbitrary reward function, possibly “dense” (i.e., per-token). Equation˜22 is known as maximum entropy RL. It is known that this objective is equivalent to solving a variational inference problem which discuss next.
@@
Then, Equation˜22 is equivalent to minimizing the KL divergence with respect to $\pi^{\star}$:
@@
Note that Equation˜24 is equivalent to the SDPO objective (Equation˜1) with implicit reward $r(y_{t}\mid x,y_{<t})=\log q(y_{t}\mid x,f,y_{<t})$ and $\lambda=1$. In this sense, SDPO can be seen as a maximum entropy RL algorithm with dense rewards constructed implicitly through the retrospective model.
@@ D.1 Learning without rich environment feedback
Qwen3-8B 35.6 59.2 27.9 58.9 57.5
+ GRPO 54.2 69.6 62.9 74.5 34.3 51.8 74.3 77.1 61.7 68.1
+ GRPO (on-policy) 54.2 69.6 62.9 74.8 30.3 49.4 73.3 75.8 61.7 68.1
+ SDPO (on-policy) 59.9 70.1 70.6 80.6 53.1 53.1 72.1 78.3 56.4 68.5
Olmo3-7B-Instruct 18.8 37.7 18.1 36.7 39.3
+ GRPO 42.7 54.3 55.3 63.3 54.2 63.8 73.8 78.1 56.4 65.0
+ GRPO (on-policy) 48.8 54.3 62.7 62.7 54.2 63.8 67.9 74.4 56.0 61.3
+ SDPO (on-policy) 59.2 76.8 60.3 71.4 56.1 58.3 75.3 79.2 57.3 62.5
@@ D.2.1 Additional Results
Figure 16: Accuracy (pass@1) for varying train batch sizes (4, 8, 16, 32) and number of rollouts (4, 8) for training SDPO and GRPO with Qwen3-8B (Yang et al., 2025) on LCBv6, $\pm$ stderr across 3 seeds. Different shades of the same color correspond to different runs.
@@
Figure 17: Average validation accuracy by model size, $\pm$ std across 3 seeds. With Qwen2.5-Instruct (Qwen et al., 2024) and Qwen3 (Yang et al., 2025) on LCBv6. Until step 65 for Qwen2.5 and until step 80 for Qwen3.
@@ E.3 User Templates
You are a helpful function-calling AI assistant. You are provided with function signatures within <functions></functions> XML tags. You may call one or more functions to assist with the user query. Output any function calls within <function_calls></function_calls> XML tags. Do not make assumptions about what values to plug into functions.
Listing 3: System prompt: Tool use
@@
Listing 4: Example user prompt: Tool use
@@ F.3 Environment Feedback
Listing 5: Example of feedback “Wrong Answer” from our code environment in case of a wrong answer, inspired by LeetCode
@@
Listing 6: Example of feedback “Memory Error” from our code environment in case of a wrong answer, inspired by LeetCode
@@
Listing 7: Example of feedback “Index Error” from our code environment in case of a wrong answer, inspired by LeetCode

### v2

@@ B.2 Trust-region Teacher
The solution to Equation˜17 can be expressed in closed form as
@@
Let $\alpha:=1/\lambda$. Then, the solution to Equation˜17 can be characterized in closed form as
@@ Appendix C Additional Related Work
Several recent works study dense (per-token) reward assignment given access to an external reward model, typically by exploiting the reward model’s internal structure (Chan et al., 2024; Cao et al., 2025b). Relatedly, Li et al. (2025b) argue that a token-level reward signal is implicit in an LLM’s logits by linking next-token prediction to offline inverse reinforcement learning, effectively yielding a training-free reward model for RL fine-tuning.
@@ C.1 SDPO as Maximum Entropy RL
where $\smash{\pi_{\theta}(y\mid x)=\prod_{t=1}^{T}\pi_{\theta}(y_{t}\mid x,y_{<t})}$ and $\smash{\mathrm{H}\left[\pi_{\theta}(\cdot\mid x)\right]=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}{}\left[-\log\pi_{\theta}(y\mid x)\right]}$ is the entropy of the policy. Here, $r(y_{t}\mid x,y_{<t})$ is an arbitrary reward function, possibly “dense” (i.e., per-token). Equation˜19 is known as maximum entropy RL. It is known that this objective is equivalent to solving a variational inference problem which discuss next.
@@
Then, Equation˜19 is equivalent to minimizing the KL divergence with respect to $\pi^{\star}$:
@@
Note that Equation˜21 is equivalent to the SDPO objective (Equation˜1) with implicit reward $r(y_{t}\mid x,y_{<t})=\log q(y_{t}\mid x,f,y_{<t})$ and $\lambda=1$. In this sense, SDPO can be seen as a maximum entropy RL algorithm with dense rewards constructed implicitly through the retrospective model.
@@ D.1 Learning without rich environment feedback
Qwen3-8B 41.2 59.2 30.8 58.9 57.5
+ GRPO 65.9 74.5 62.9 74.5 35.1 59.9 74.3 77.1 61.7 68.1
+ GRPO (on-policy) 52.2 71.6 62.9 74.8 49.8 49.8 73.3 75.8 61.7 68.1
+ SDPO (on-policy) 73.2 80.9 70.6 80.6 50.6 56.8 72.1 78.3 56.4 68.5
Olmo3-7B-Instruct 22.8 37.7 16.2 36.7 39.3
+ GRPO 53.1 67.7 55.3 63.3 35.6 55.8 73.8 78.1 56.4 65.0
+ GRPO (on-policy) 47.1 65.4 62.7 62.7 49.8 49.8 67.9 74.4 56.0 61.3
+ SDPO (on-policy) 68.0 80.0 60.3 71.4 48.0 52.8 75.3 79.2 57.3 62.5
@@ D.2.1 Additional Results
Figure 16: Accuracy (pass@1) for varying train batch sizes (4, 8, 16, 32) and number of rollouts (4, 8) for training SDPO and GRPO with Qwen3-8B (Yang et al., 2025a) on LCBv6, $\pm$ stderr across 3 seeds. Different shades of the same color correspond to different runs.
@@
Figure 17: Average validation accuracy by model size, $\pm$ std across 3 seeds. With Qwen2.5-Instruct (Qwen et al., 2024) and Qwen3 (Yang et al., 2025a) on LCBv6. Until step 65 for Qwen2.5 and until step 80 for Qwen3.
@@ E.3 User Templates
Listing 3: Example user prompt: Tool use
@@ F.3 Environment Feedback
Listing 4: Example of feedback “Wrong Answer” from our code environment in case of a wrong answer, inspired by LeetCode
@@
Listing 5: Example of feedback “Memory Error” from our code environment in case of a wrong answer, inspired by LeetCode
@@
Listing 6: Example of feedback “Index Error” from our code environment in case of a wrong answer, inspired by LeetCode